Scores on the math portion of the SAT (SAT-M) in a recent year have followed a normal distribution with mean μ = 507 and standard deviation σ = 111.

What is the probability that the mean SAT-M score of a random sample of 4 students who took the test that year is more than 600?

If normality conditions are met, round your Z-score calculation to 2 decimal places.

Normal Distribution Calculator (Standard Normal)

Group of answer choices

1. P(Z>1.68) = 0.05

2. P(Z>0.84)=0.20P(Z>0.84)=0.20P(Z>0.84)=0.20P(Z>0.84)=0.20

3. It is impossible to tell because normality conditions are not met
Answer #1

It is given that scores on the math portion of the SAT (SAT-M) in a recent year have followed a normal distribution with mean μ = 507 and standard deviation σ = 111.

Further, let 1. , X2, X3, X4 denote the SAT-M score of a random sample of 4 students. Then X; N (507,1112); i = 1,2,3,4

Now we know that if \small X_1, \dots, X_n is a random sample from \small \mathcal{N}(\mu, \sigma^2) distribution, then the sample mean \small \bar{X} also follows Normal distribution with the same mean but a scaled variance, that is \small \bar{X} \sim \mathcal{N}(\mu, \frac{\sigma^2}{n}) .

Thus, if 1 X = (X1 + X2 + X3 + X4) is the mean score of the random sample of 4 students, then, 111? XN (507, . Hence, the normality assumption holds.

Now, we want to calculate P(X > 600 . Towards that, we calculate the z-score as,

\small z = \frac{X - \mu}{\sigma} = \frac{600 - 507}{\sqrt{\frac{111^2}{4}}} = \frac{93*2}{111} = 1.6757

Rounding off to 2-decimal places, we get z = 1.68. Thus, the required probability is

\small P(\bar{X} > 600) = P(Z > z) = P(Z > 1.68 )
Hence, option (A) is correct.

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

111? XN (507,

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image